OpenStudy (anonymous):

How does x/(x+1) = 1-1/(x+1)

OpenStudy (anonymous):

Do you know about LCM or LCD ???

OpenStudy (lgbasallote):

you can split up the numerator \[\implies \frac{x+1 - 1}{x+1}\] now split up the fraction \[\implies \frac{x+1}{x+1} - \frac{1}{x+1}\] does that give you a hint?

OpenStudy (chihiroasleaf):